If 2.09g of the vapor of a volatile liquid is able to fill a 498-mL flask at 370C and 617mm Hg, what is the molecular weight of the liquid?
Answer #1

PV = nRT

P = 0.812 atm

V = 0.498 L

T = 310K

0.812*0.498 = n*0.0821*310

0.404 = n * 25.5

n = 0.404/25.5 = 0.0158 mole

0.0158 mole = 2.09 g / molar mass

molarn mass = 2.09/0.0158 = 132.3 g. mol^-1